Solve for c1 and c2 for the case of the two-period consumption and saving model with certainty, using the quadratic form u(c) = c - 0.5ac2 for the instantaneous utility function and with the following income patterns {y1, y2} and initial wealth w0.


1
Expert's answer
2021-09-23T14:40:39-0400

Consider

"u(c) = c-0.5ac^2 \\

c_1+b_1=y_1 \\

c_2+b_2 = y_2 + (1+ω)b_1 \\

c_1 + \frac{c_2}{1+ω} = y_1 + \frac{y_2}{1+ω}"

Consider

"c_1 + \frac{c_2}{1+ω}=x"

Define

"u'(c_1)dc_1 + βu'(c_2)dc_2 = 0 \\

\frac{dc_2}{dc_1} = \frac{-u'(c_1)}{β(u')c_2} \\

u(c_1>c_2) = c_1 -0.5ac_1^2 \\

= 1 -ac_1"

Similary "(1-ac_2)"

Consider "c_1=c_2"

"\frac{c(1-a)}{c(1-a)}=1"

However

"\frac{dc_2}{dc_1} = \frac{-c_2}{βc_1} = \frac{1}{β}"